# Economic DAO Governance: A Contestable Control Approach

### v3

@@ 5.4 Post-auction Incentives and EV-robustness
It is clear that no arrangement is going to be EV-robust with respect to token holdings. It is easy enough to hedge them in a way that is equivalent to a sale. But the same cannot be said of the value deposit and security deposit, which are key deterrents to value destruction and simultaneously incentivize achievement of the business plan. These deposits are EV-robust. Hedging the value deposit requires purchasing the opposite derivatives position, which in the case of the standard value deposit forfeit function effectively means buying a put with a $V^{*}$ strike price and selling a put with a $P_{0}$ strike price. [^76] See note 32 supra. This position would be expensive in the sense that it would cost a significant proportion of the deposit. With respect to outcomes below $P_{0}$, a large put position with appropriate strike prices would be required to offset potential forfeits involving both the value deposit and surety deposit. This put position also would be costly and possibly difficult to assemble without moving the token price significantly downward while assembling the position.
@@ 5.6 Decentralization, Regulation, and Legal Aspects
A first approach is decentralization-focused in the sense that the degree of decentralization determines whether and to what extent DAOs and other cryptocurrency applications are regulated. There are many variants that fall into this category. Under the recently implemented European Union laws governing cryptocurrencies, the key role of decentralization is explicit with an exemption from regulation for the case in which “crypto-asset services … are provided in a fully decentralized manner without any intermediary.” [^81] European Parliament and the Council of the European Union (2023) (Directive (22)). In the United States, the test set forth by the Supreme Court in SEC v W.J. Howey Co.828282328 U.S. 293 (1946). determines which cryptocurrencies will be considered “securities,” subject to registration with the Securities and Exchange Commission (“SEC”) and on-going regulation by the Commission. Cryptocurrencies that are not “securities” generally fall under the regulatory ambit of the Commodities Futures Trading Commission (“CFTC”) as “commodities.” [^83] A cryptocurrency can be both a “security” and a “commodity” under the applicable statures and regulations, creating joint regulatory oversight by the SEC and CFTC. Virtually all cryptocurrencies are considered to be ”commodities,” and the CFTC has regulatory authority to police fraud and market manipulation in spot markets for commodities. The SEC has the same authority with respect to cryptocurrencies that also are “securities.”
One branch of the Howey test requires investor dependence on the “efforts of others” to make a profit as a necessary condition for a token to be a security.848484328 U.S. 293, 299 (1946). If the token is decentralized, in the extreme case just code running on its own, arguably there is no such dependence. Furthermore, disclosure of the backgrounds and plans of the key managing parties is a major objective of registration and on-going regulation under the securities laws. [^85] Hinman (2018). If there are no such parties, regulation loses its primary rationale. In addition, if there are no managing parties, it is unclear who would be available to comply with regulation in any event.
@@
The SEC went on to discuss compliance costs. The SEC noted that “factors associated with certain technologies … might increase compliance costs” and that compliance might ”significantly reduce the extent to which the system is ‘decentralized’ or otherwise operates in a manner consistent with the principles that the crypto asset industry commonly refer to as ‘DeFi.’” [^99] Id. at 122. The SEC singled out DAOs as a case for which the highest possible compliance costs might occur.100100100The Commission put it this way (Id. at 124-125): The Commission preliminarily believes that a reasonable case, in which the highest possible compliance costs would result, would be a … system that performs exchange activities in part using smart contracts, but in which control over changes to the smart contracts is given to a token-based voting mechanism, which may use governance tokens as discussed above, and where the tokens are dispersed among a large number of investors. The SEC continued, stating that “the holders of governance tokens, or other tokens that carry voting rights, may bear the responsibility of ensuring compliance with the system,” suggesting that the token holders should create an organization or delegate persons to undertake the activities required for compliance, putting money into the DAO if the costs are not covered by fees.101101101Id. Furthermore, the token holders would have to have the power to alter the relevant smart contracts by vote if necessary for compliance, and in the case of immutable smart contracts, miners or validators would have to alter the smart contracts via a hard fork.102102102Id. at 126-127.
@@
While it is true that these positions were asserted in a Release soliciting additional comments, they are the responses to concerns raised by previous comments suggesting that the scope of the Rules should be limited to create an exemption or at least breathing space for decentralized projects, including DAOs. And although the Rules only apply to exchanges, the same reasoning would apply to inclusion of DAOs in the general category of “securities” subject to regulation.103103103This reasoning would greatly weaken the scope of the “efforts of others” branch of Howey as a restrictive necessary condition for a token to be considered a security, and, if pursued, most likely would trigger legal challenges and a possible reconsideration or clarification of Howey itself.
@@
It also is noteworthy that the SEC was deeply divided concerning the position taken in the Reopening Release, with the minority group in the 3-2 Commission split being sharply critical of that position. Commissioner Hester Peirce articulated the critique in a dissent to the Reopening Release.104104104Peirce (2023). She characterized the Release position as an “approach to exchange regulation as something that must not – indeed cannot – be altered to allow room for new technologies or for new ways of doing business.”105105105Id. In a later videotaped interview, Commissioner Peirce took the position that the Commission should “set true decentralized protocols aside” because the “technology itself is a substitute for many of the regulatory protections” otherwise required when interactions are with “an intermediary, a centralized firm.”106106106Hamilton (2023). Choi et al. (2021, pp. 2-3, 7-9) describe this situation as one that involves “functional and regulatory equivalence.” The application is the same as one that normally would be regulated, functional equivalence, and elements in the technology achieve the same purpose as regulation, regulatory equivalence. She noted that you can lose money and get hurt but “you are opting into that when you are entering into decentralized transactions.”107107107Id. This approach is very similar in flavor to the newly enacted E.U. rule that exempts a crypto application that operates “in a fully decentralized manner without any intermediary.”
@@
A third regulatory approach, which can be termed regulatory gradualism, consists of observing a new technology and its operation and then, after a period of observation, deciding on a new, appropriate regulatory regime. Part of Commissioner Peirce’s dissent to the Reopening Release advocated this approach. She pointed out with approbation that the SEC had used a regulatory gradualism approach in the 1990s, permitting new, electronic exchanges to operate for a period of a few years despite being non-compliant with the prevailing strict regulatory requirements for exchanges, resulting in successful development and implementation of a new body of regulations tailored to such exchanges. Closely related to this 1990s example is the idea of creating a “regulatory sandbox” in which new cryptocurrency technologies operate and within which participants understand that they are on their own with respect to possible losses prior to the imposition of an existing regulatory regime or the implementation of a new one.108108108Commissioner Peirce has advocated this idea separately from her dissent from the Reopening Release. See Piro (2021). If the dissenters gain a majority on the Commission in the future, policy may well move toward this third approach or back to the first decentralization-focused approach that would follow from classical adherence to Howey.109109109The SEC has five Commissioners appointed by the President, but only three can be from the same party. The current 3-2 split is along party lines, with the Republican Commissioners in the minority. Thus, the regulatory approach can change sharply every time there is a Presidential election.
@@
Use of the mechanism raises four sets of regulatory considerations. First, there are the potential impacts of shifting to the mechanism under existing regulation. Second, the mechanism creates regulatory opportunities, ways in which DAOs could more easily fit within regulatory regimes without adverse impacts on their mission and structure. Third, regulation may directly impact the effectiveness of the mechanism. Fourth, the mechanism raises regulatory questions if the DAO is subject to the core aspects of securities or commodities regulation.
@@
The mechanism creates some regulatory opportunities, ways in which a DAO may fit more easily within regulatory regimes without compromising its central purposes. First, at least during the control period, there is clear party who can coordinate and interact with regulators, parallel to the managers of a traditional corporation. That is not the case if there is only implicit control or a group of large token block holders. The ability of the control party to interact with regulators creates a shield for other parties such as developers, validators, and token holders. This shield may be especially valuable in the face of a joint and several responsibility regulatory approach such as the one suggested by the SEC in the Release that envisions a very broad set of actors being subject to potential regulatory obligations and liabilities. Otherwise, this approach might well, in the words of dissenting Commissioner Peirce, “render innovation kaput.”110110110Peirce (2023).
Second, various aspects of the mechanism create a case for establishing an exemptive safe harbor or reduced regulatory regime for DAOs that use it. As we have seen, there is a strong argument that the mechanism enhances decentralization in the face of both implicit and explicit control possibilities, and in many cases actualities, that emerge from token ownership and voting regimes. The mechanism also aims at the socially optimal outcome in each auction. And during control periods, there is a party that is an ideal subject of disclosure, the central goal of securities regulation. All of these qualities make a DAO using the mechanism a very attractive candidate for inclusion in a regulatory sandbox or for temporary forbearance by regulators under regulatory gradualism approaches.
@@
Some aspects of regulation may make it easier to implement the mechanism itself. As discussed above, one of the biggest challenges for the mechanism is toehold reporting. If the DAO is a security, then Schedule 13D reporting or a modified version of it would apply, creating a legal obligation to report a non-negligible toehold position.111111111Under section 13(d) of the Securities Act of 1934 any person who “directly or indirectly” becomes the beneficial owner of more than 5% of any class of registered equity securities, and certain equity securities that are exempt from registration must, within ten days, report certain information to the issuer, to any exchange on which the security is traded, and to the SEC. “[I]f the purpose of the purchases or prospective purchases is to acquire control of the business of the issuer of the securities,” the information must include “any plans or proposals which such persons may have to liquidate such issuer, to sell its assets to or merge it with any other persons, or to make any other major change in its business or corporate structure. Securities Exchange Act of 1934, Pub. L. No. 73-291, 48 Stat. 881 (codified at 15 U.S.C. §78a (2012)). Schedule 13D is the form used to make the disclosures required by section 13(d). This legal obligation and its enforcement is external, not a code feasible part of the mechanism itself. Nonetheless, it could go a long way toward filling what might be a significant implementation gap for the mechanism.
@@
Even if the DAO is not considered a security under U.S. law, it will be considered a commodity if it is publicly traded and thus subject to the ability of the CFTC to police fraud. The DAO can leverage this potential fraud liability with respect to the toehold position. Being able to submit an artificially low bid by failing to report the entire toehold position is fraud. The full enforcement and adjudication power of the CFTC comes into play. The control party would be subject to discovery and required testimony under oath. Hiding part of the toehold would create potentially serious consequences for the control party. This potential fraud liability is another external device, not part of the code feasible core of the DAO. But again, it can help address a difficult implementation gap for the DAO.112112112It also might be possible to create some structure in the DAO to aid the applicability of this liability. For instance, the DAO might trigger notice of an auction and the toehold reporting obligation to the CFTC, or, if the DAO is a security, to the SEC, along with disclosure of the identity of the control party. The CFTC and the SEC could incorporate this notice into their regulatory regimes as a trigger for possible inquiry.
@@
Finally, there are legal considerations in addition to regulation. Most prominent is the issue of legal liability for DAO participants. This issue exists with or without the mechanism. DAOs can be “wrapped” in LLC or other legal forms to provide limited liability for participants, including token holders and any parties that might be seen as “management.”113113113See, e.g., Guillaume and Riva (2022a), Guilluame and Riva (2022b), Kerr and Jennings (2022), Jennings and Kerr (2022), and Brummer and Seira (2022). Limited liability might be particularly valuable for the control party due to the high visibility of being in that position. That is the main new element created by the mechanism. The question of the legal status and use of wrappers is a quickly developing field and the choices are complex. We do not discuss that question further here.

### v4

@@ 5.4 Post-auction Incentives and EV-robustness
It is clear that no arrangement is going to be EV-robust with respect to token holdings. It is easy enough to hedge them in a way that is equivalent to a sale. But the same cannot be said of the value deposit and security deposit, which are key deterrents to value destruction and simultaneously incentivize achievement of the business plan. These deposits are EV-robust. Hedging the value deposit requires purchasing the opposite derivatives position, which in the case of the standard value deposit forfeit function effectively means buying a put with a $V^{*}$ strike price and selling a put with a $P_{0}$ strike price. [^75] See note 31 supra. This position would be expensive in the sense that it would cost a significant proportion of the deposit. With respect to outcomes below $P_{0}$, a large put position with appropriate strike prices would be required to offset potential forfeits involving both the value deposit and surety deposit. This put position also would be costly and possibly difficult to assemble without moving the token price significantly downward while assembling the position.
@@ 5.6 Decentralization, Regulation, and Legal Aspects
After introducing and discussing the three regulatory approaches, we turn to the interaction of the mechanism with regulation. We begin with the most important and general point, that by providing additional protection for investors, adding the mechanism greatly strengthens the case for a much less burdensome regulatory regime which at the same time is likely to be more effective. Following that general discussion we consider ways in which the mechanism interacts with existing regulation under the three approaches.
5.6.1 Three Regulatory Approaches
A first approach is decentralization-focused in the sense that the degree of decentralization determines whether and to what extent DAOs and other cryptocurrency applications are regulated. There are many variants that fall into this category. Under the recently implemented European Union laws governing cryptocurrencies, the key role of decentralization is explicit with an exemption from regulation for the case in which “crypto-asset services … are provided in a fully decentralized manner without any intermediary.” [^80] European Parliament and the Council of the European Union (2023) (Directive (22)). In the United States, the test set forth by the Supreme Court in SEC v W.J. Howey Co.818181328 U.S. 293 (1946). determines which cryptocurrencies will be considered “securities,” subject to registration with the Securities and Exchange Commission (“SEC”) and on-going regulation by the Commission. Cryptocurrencies that are not “securities” generally fall under the regulatory ambit of the Commodities Futures Trading Commission (“CFTC”) as “commodities.” [^82] A cryptocurrency can be both a “security” and a “commodity” under the applicable statures and regulations, creating joint regulatory oversight by the SEC and CFTC. Virtually all cryptocurrencies are considered to be ”commodities,” and the CFTC has regulatory authority to police fraud and market manipulation in spot markets for commodities. The SEC has the same authority with respect to cryptocurrencies that also are “securities.”
One branch of the Howey test requires investor dependence on the “efforts of others” to make a profit as a necessary condition for a token to be a security.838383328 U.S. 293, 299 (1946). If the token is decentralized, in the extreme case just code running on its own, arguably there is no such dependence. Furthermore, disclosure of the backgrounds and plans of the key managing parties is a major objective of registration and on-going regulation under the securities laws. [^84] Hinman (2018). If there are no such parties, regulation loses its primary rationale. In addition, if there are no managing parties, it is unclear who would be available to comply with regulation in any event.
@@ 5.6.1 Three Regulatory Approaches
The SEC went on to discuss compliance costs. The SEC noted that “factors associated with certain technologies … might increase compliance costs” and that compliance might ”significantly reduce the extent to which the system is ‘decentralized’ or otherwise operates in a manner consistent with the principles that the crypto asset industry commonly refer to as ‘DeFi.’” [^98] Id. at 122. The SEC singled out DAOs as a case for which the highest possible compliance costs might occur. [^99] The Commission put it this way (Id. at 124-125): The Commission preliminarily believes that a reasonable case, in which the highest possible compliance costs would result, would be a … system that performs exchange activities in part using smart contracts, but in which control over changes to the smart contracts is given to a token-based voting mechanism, which may use governance tokens as discussed above, and where the tokens are dispersed among a large number of investors. The SEC continued, stating that “the holders of governance tokens, or other tokens that carry voting rights, may bear the responsibility of ensuring compliance with the system,” suggesting that the token holders should create an organization or delegate persons to undertake the activities required for compliance, putting money into the DAO if the costs are not covered by fees.100100100Id. Furthermore, the token holders would have to have the power to alter the relevant smart contracts by vote if necessary for compliance, and in the case of immutable smart contracts, miners or validators would have to alter the smart contracts via a hard fork.101101101Id. at 126-127.
@@
While it is true that these positions were asserted in a Release soliciting additional comments, they are the responses to concerns raised by previous comments suggesting that the scope of the Rules should be limited to create an exemption or at least breathing space for decentralized projects, including DAOs. And although the Rules only apply to exchanges, the same reasoning would apply to inclusion of DAOs in the general category of “securities” subject to regulation.102102102This reasoning would greatly weaken the scope of the “efforts of others” branch of Howey as a restrictive necessary condition for a token to be considered a security, and, if pursued, most likely would trigger legal challenges and a possible reconsideration or clarification of Howey itself.
@@
It also is noteworthy that the SEC was deeply divided concerning the position taken in the Reopening Release, with the minority group in the 3-2 Commission split being sharply critical of that position. Commissioner Hester Peirce articulated the critique in a dissent to the Reopening Release.103103103Peirce (2023). She characterized the Release position as an “approach to exchange regulation as something that must not – indeed cannot – be altered to allow room for new technologies or for new ways of doing business.”104104104Id. In a later videotaped interview, Commissioner Peirce took the position that the Commission should “set true decentralized protocols aside” because the “technology itself is a substitute for many of the regulatory protections” otherwise required when interactions are with “an intermediary, a centralized firm.”105105105Hamilton (2023). Choi et al. (2021, pp. 2-3, 7-9) describe this situation as one that involves “functional and regulatory equivalence.” The application is the same as one that normally would be regulated, functional equivalence, and elements in the technology achieve the same purpose as regulation, regulatory equivalence. She noted that you can lose money and get hurt but “you are opting into that when you are entering into decentralized transactions.”106106106Id. This approach is very similar in flavor to the newly enacted E.U. rule that exempts a crypto application that operates “in a fully decentralized manner without any intermediary.”
@@
A third regulatory approach, which can be termed regulatory gradualism, consists of observing a new technology and its operation and then, after a period of observation, deciding on a new, appropriate regulatory regime. Part of Commissioner Peirce’s dissent to the Reopening Release advocated this approach. She pointed out with approbation that the SEC had used a regulatory gradualism approach in the 1990s, permitting new, electronic exchanges to operate for a period of a few years despite being non-compliant with the prevailing strict regulatory requirements for exchanges, resulting in successful development and implementation of a new body of regulations tailored to such exchanges. Closely related to this 1990s example is the idea of creating a “regulatory sandbox” in which new cryptocurrency technologies operate and within which participants understand that they are on their own with respect to possible losses prior to the imposition of an existing regulatory regime or the implementation of a new one.107107107Commissioner Peirce has advocated this idea separately from her dissent from the Reopening Release. See Piro (2021). If the dissenters gain a majority on the Commission in the future, policy may well move toward this third approach or back to the first decentralization-focused approach that would follow from classical adherence to Howey.108108108The SEC has five Commissioners appointed by the President, but only three can be from the same party. The current 3-2 split is along party lines, with the Republican Commissioners in the minority. Thus, the regulatory approach can change sharply every time there is a Presidential election.
5.6.2 Interaction of the Mechanism with Regulation
A key issue for both DAOs in general and for the mechanism in particular is the extent to which the associated technologies reduce the need for regulation. “Regulatory equivalence” exists if elements in the technology achieve the same purpose as regulation.109109109See supra note 105. For example, to the extent DAOs and blockchain-based applications replace intermediaries with code, they eliminate the need for regulation to address potential misbehavior by such intermediaries.
Blockchain elements that typically accompany DAOs make not only the governing code public and transparent but also much if not all of the operations conducted by the DAO. This transparency reduces the need for many of the disclosures that regulators typically require in order to protect investors by giving them information about their investments. Requiring such disclosures is a major focus of regulation in the United States and many other jurisdictions.110110110The preamble of the Securities Act of 1933, one of the two key statutes that form a basis for securities regulation in the United States, describes the goals of the statute as follows: “AN ACT To provide full and fair disclosure of the character of securities sold in interstate and foreign commerce and through the mails, and to prevent frauds in the sale thereof, and for other purposes.” Securities Act of 1933, 15 U.S.C. §§ 77a-77aa (2018).
In addition to requiring disclosure, regulatory systems typically include market structure and procedural protections for investors, often aimed at promoting fairness for such investors as they interact with sophisticated parties or exchanges.111111111A good statement of these purposes is the preamble of the Securities Exchange Act of 1934, the second in time of the two key acts aimed at protecting investors in the United States: “AN ACT To provide for the regulation of securities exchanges and of over-the-counter markets operating in interstate and foreign commerce and through the mails, to prevent inequitable and unfair practices on such exchanges and markets, and for other purposes.” Securities Exchange Act of 1934 (June 6, 1934, ch. 404, title I, § 1, 48 Stat. 881). An example is the rules governing tender offers in the United States, offers that a bidder makes to shareholders of public companies to buy their shares for a particular price at a specified time. These rules require bidders to leave any offering open for at least 20 days, to refrain from giving different subscribers different terms, and to make certain disclosures.11211211215 U.S.C. § 78m; 17 CFR §240.13e, §240.14d, and §240.14e Furthermore, management of the target corporation is required to issue its opinion of the offering terms.113113113Rule 14d-9 [17 CFR §240.14d-9] and Schedule 14D-9 [17 CFR §240.14d-101].
The mechanism goes considerably beyond these customary disclosure-based, market structure, and procedural protections to provide direct substantive protection to investors. Consider, first, control period situations. Holding tokens at the beginning of an auction protects holders against any losses during the control period and guarantees that they will realize the increase in value or its financial equivalent promised by the control party on the portion of the tokens that they retain at the end of the auction. This arrangement exists whether or not the auction results in a shift in control as opposed to situations in which the winning bidder previously had control via an earlier auction or from a token position combined with active participation in governance during an open period. More generally, the auction aims to transfer as much of the social surplus from the control party’s operations as possible to these token holders. This high degree of substantive protection in the face of shifts in control does not exist under current regulatory regimes.
The control period protections are not operative for token holders during open periods and do not apply to tokens purchased during a control period after the auction has been completed.114114114The protection against losses and the guarantee of claimed gains applies to the particular holders at the time an auction is completed rather than attaching to the tokens themselves. Thus, buyers during a control period do not enjoy these protections. See supra page 4.3.1. Passive, “buy and hold” investors enjoy particularly strong benefits overall because they always benefit from the control period protections. But token holders in both situations enjoy two other protections. First, the threat of future auctions has a disciplinary effect with respect to misbehavior or poor performance on the part of existing governing parties, including the case in which a mass of diffuse token holders determine the DAO’s path.115115115Because auctions may be freely initiated at any time and because the auctions are not subject to the restrictions that typically apply in corporate regulatory regimes, the threat implicit from auctions under the mechanism is most likely a much more effective external governance device than the threat from conventional corporate takeovers. Second, these groups will enjoy future control period protections if subsequent auctions arise.
The substantive investor protections inherent in the mechanism both substantially strengthen the case for applying a less burdensome regulatory regime to DAOs that employ it and create the prospect of more fully protecting investors than existing regulatory regimes. They also make a DAO using the mechanism a very attractive candidate for inclusion in a regulatory sandbox or for temporary forbearance by regulators under regulatory gradualism approaches.
There are additional ways in which the mechanism advances the case for a milder form of regulation as well as for regulatory gradualism. As we have seen, there is a strong argument that the mechanism enhances decentralization in the face of both implicit and explicit control possibilities, and in many cases actualities, that emerge from token ownership and voting regimes. The mechanism induces choice of the path forward that will maximize social surplus and attempts to distribute that surplus in a manner that allows projects to reflect their true social value for purposes of initial and on-going investment. Excessive regulation would imperil these benefits. Finally, disclosure of the identity of the guiding parties of an enterprise is a major goal of securities regulation. The diffusion of control in DAOs tends to make identification of managing parties and elucidation of information about them difficult if not impossible. In contrast, by the operation of the mechanism, the identity of the managing parties will be evident, at least during control periods.
Do all of these features add up to a case for exemption? The answer to that question is not clear. There may remain some scope for regulation. In particular, disclosure of certain information beyond that available from the transparency inherent in blockchain aspects may have value to market participants and potential auction bidders. We leave exploration of that possibility and the question of exemption to future work.
@@ 5.6.2 Interaction of the Mechanism with Regulation
Aside from creating a general case for reduced regulation or a regulatory gradualism approach, use of the mechanism raises four additional considerations with respect to existing regulation under the three approaches. First, there are the potential regulatory impacts from introducing the mechanism into DAO governance systems. Second, the mechanism creates regulatory opportunities, ways in which DAOs could more easily fit within existing regulatory regimes without adverse impacts on their mission and structure. Third, regulation may directly impact the effectiveness of the mechanism. Fourth, the mechanism raises regulatory questions if the DAO is subject to the core aspects of securities or commodities regulation.
@@
The mechanism creates some regulatory opportunities, ways in which a DAO may fit more easily within regulatory regimes without compromising its central purposes. At least during control periods, there is clear party who can coordinate and interact with regulators, parallel to the managers of a traditional corporation. That is not the case if there is only implicit control or a group of large token block holders. The ability of the control party to interact with regulators creates a shield for other parties such as developers, validators, and token holders. This shield may be especially valuable in the face of a joint and several responsibility regulatory approach such as the one suggested by the SEC in the Release that envisions a very broad set of actors being subject to potential regulatory obligations and liabilities. Otherwise, this approach might well, in the words of dissenting Commissioner Peirce, “render innovation kaput.”116116116Peirce (2023).
@@
Some aspects of regulation may make it easier to implement the mechanism itself. As discussed above, one of the biggest challenges for the mechanism is toehold reporting. If the DAO is a security, then Schedule 13D reporting or a modified version of it would apply, creating a legal obligation to report a non-negligible toehold position.117117117Under section 13(d) of the Securities Act of 1934 any person who “directly or indirectly” becomes the beneficial owner of more than 5% of any class of registered equity securities, and certain equity securities that are exempt from registration must, within ten days, report certain information to the issuer, to any exchange on which the security is traded, and to the SEC. “[I]f the purpose of the purchases or prospective purchases is to acquire control of the business of the issuer of the securities,” the information must include “any plans or proposals which such persons may have to liquidate such issuer, to sell its assets to or merge it with any other persons, or to make any other major change in its business or corporate structure.” Securities Exchange Act of 1934, Pub. L. No. 73-291, 48 Stat. 881 (codified at 15 U.S.C. §78a (2012)). Schedule 13D is the form used to make the disclosures required by section 13(d). This legal obligation and its enforcement is external, not a code feasible part of the mechanism itself. Nonetheless, it could go a long way toward filling what might be a significant implementation gap for the mechanism.
@@
Even if the DAO is not considered a security under U.S. law, it will be considered a commodity if it is publicly traded and thus subject to the ability of the CFTC to police fraud. The DAO can leverage this potential fraud liability with respect to the toehold position. Being able to submit an artificially low bid by failing to report the entire toehold position is fraud. The full enforcement and adjudication power of the CFTC comes into play. The control party would be subject to discovery and required testimony under oath. Hiding part of the toehold would create potentially serious consequences for the control party. This potential fraud liability is another external device, not part of the code feasible core of the DAO. But again, it can help address a difficult implementation gap for the DAO.118118118It also might be possible to create some structure in the DAO to aid the applicability of this liability. For instance, the DAO might trigger notice of an auction and the toehold reporting obligation to the CFTC, or, if the DAO is a security, to the SEC, along with disclosure of the identity of the control party. The CFTC and the SEC could incorporate this notice into their regulatory regimes as a trigger for possible inquiry.
@@
Finally, there are legal considerations in addition to regulation. Most prominent is the issue of legal liability for DAO participants. This issue exists with or without the mechanism. DAOs can be “wrapped” in LLC or other legal forms to provide limited liability for participants, including token holders and any parties that might be seen as “management.”119119119See, e.g., Guillaume and Riva (2022a), Guilluame and Riva (2022b), Kerr and Jennings (2022), Jennings and Kerr (2022), and Brummer and Seira (2022). Limited liability might be particularly valuable for the control party due to the high visibility of being in that position. That is the main new element created by the mechanism. The question of the legal status and use of wrappers is a quickly developing field and the choices are complex. We do not discuss that question further here.